OpenStudy (anonymous):

42 is 7% of what number

OpenStudy (anonymous):

A MOD IS VIEWING THIS? O_O

OpenStudy (phi):

42 is 7% of x 42 = 0.07 x solve for x by dividing both sides by 0.07 42/0.07 = (0.07/0.07) x or x= 42/0.07

OpenStudy (phi):

if you write 0.07 as 7/100 then you would have \[ \frac{7}{100} x = 42 \] multiply both sides by 100/7 \[ \frac{100}{7} \cdot \frac{7}{100} x = 42\cdot \frac{100}{7} \\ x= \frac{42}{7} \cdot 100 \]

OpenStudy (phi):

either way gives the same number.

OpenStudy (anonymous):

thanks